### v2

# Generative AI for Physical-Layer Authentication

Abstract
In recent years, Artificial Intelligence (AI)-driven Physical-Layer Authentication (PLA), which focuses on achieving endogenous security and intelligent identity authentication, has attracted considerable interest. When compared with Discriminative AI (DAI), Generative AI (GAI) offers several advantages, such as fingerprint data augmentation, fingerprint denoising and reconstruction, and protection against adversarial attacks. Inspired by these innovations, this paper provides a systematic exploration of GAI’s integration into PLA frameworks. We commence with a review of representative authentication techniques, emphasizing PLA’s inherent strengths. Following this, we revisit four typical GAI models and contrast the limitations of DAI with the potential of GAI in addressing PLA challenges, including insufficient fingerprint data, environment noises and inferences, perturbations in fingerprint data, and complex tasks. Specifically, we delve into providing GAI-enhanced methods for PLA across the fingerprint collection, model training, and performance optimization phases in detail. Moreover, we present a case study that combines fingerprint extrapolation and Generative Diffusion Model (GDM) to illustrate the superiority of GAI in bolstering the reliability of PLA. Additionally, we outline potential future research directions for GAI-based PLA.
Index Terms:
Generative AI, physical-layer authentication, 6G, identity security.
I Introduction
After 5G technology has fully entered the commercialization stage, research on 6G standards has commenced. According to the IMT-2030 framework issued by the International Telecommunication Union (ITU), 6G not only enhances the three original fundamental scenarios of 5G but also introduces three new scenarios: AI and communication, ubiquitous connectivity, and Integrated Sensing and Communication (ISAC). However, as 6G drives the Internet of everything, it may potentially introduce complex identity security threats.
Physical-Layer Authentication (PLA) technology, with its endogenous security mechanisms and lightweight characteristics, has been considered a promising method to compensate for the shortcomings of traditional secure access mechanisms[1, 2, 3]. It is primarily implemented based on physical-layer fingerprints, including channel fingerprints [3, 4] and radio frequency (RF) fingerprints [5, 6]. Traditionally, PLA is modeled as a hypothesis testing process. In recent years, AI-based PLA has gradually become a research hotspot [7].
Traditional Discriminative AI (DAI) utilizes deep neural networks to learn features from large-scale fingerprint datasets. However, these schemes rely on extensively labeled fingerprints for training. Fortunately, the emergence of Generative AI (GAI) has mitigated these limitations. Instead of learning class boundaries, GAI learns the probability distribution of training fingerprints to generate new fingerprint samples [8], which is expected to be applied to Internet of Things (IoT) [1, 9], Industrial IoT (IIoT) [3], Internet of Vehicles (IoV) [10], and satellite communication systems [6]. Nevertheless, research on GAI-based PLA is still in its early phases. The full potential of applying GAI to PLA has yet to be comprehensively investigated. Specifically, the following questions remain unanswered:
  • Q1: Why is GAI suitable for PLA?
  • Q2: What issues can GAI handle in the context of PLA?
  • Q3: How does GAI handle these issues?
To address these challenges, we review representative identity authentication approaches and GAI models, and assess the feasibility of applying GAI to PLA. The primary contributions are outlined as:
  • Firstly, we revisit classic identity authentication techniques and prominent GAI models. Then, we compare the limitations of DAI with the potential of GAI in PLA.
  • We sort out that GAI can solve the issues of PLA from the perspectives of fingerprint collection, model training, and performance optimization.
  • We present a case study to demonstrate the effectiveness of GAI in enhancing authentication performance. Additionally, we envision future research directions for GAI-based PLA.
II Overview of Identity Authentication and Generative Artificial Intelligence (GAI)
II-A Identity Authentication
II-A1 Why Identity Authentication?
Identity authentication serves as a critical component in ensuring wireless identity security, and its necessity is primarily reflected as follows.
  • Preventing Unauthorized Access: The open nature of wireless channels exposes signals to public environments, allowing attackers to impersonate legitimate terminals, steal sensitive information, and even tamper with data streams through Man-in-the-Middle attacks. Identity authentication verifies the authenticity of communicating parties, blocks unauthorized devices at the source, and establishes a trusted communication link [3].
  • Safeguarding Data Security: Enabling encryption without prior authentication may allow attackers to use forged identities to participate in key negotiation processes, resulting in encryption key leakage. Identity authentication acts as a trust anchor for encryption systems, ensuring that keys are exclusively held by legitimate parties, thereby preserving the confidentiality and integrity of data transmission [9].
  • Mitigating Resource Abuse: Identity authentication can reject devices with forged identities from occupying channel resources or launching Distributed Denial-of-Service (DDoS) attacks, thereby preventing network congestion and enhancing Quality of Service (QoS) [7].
II-A2 How to Identity Authentication?
Cryptography-based authentication is primarily implemented through cryptographic algorithms in the upper-layers of communication systems; Blockchain-based authentication represents a distributed architecture-based group identification authentication method, with the characteristics of a multi-layer framework and a fault-tolerant mechanism; Physical-Layer Key-based Authentication (PLKA) is realized using physical-layer key, which is generated based on the wireless channel reciprocity; PLA is realized based on physical-layer attributes. The detailed descriptions are shown in Table I.
TABLE I: Descriptions of four representative identity authentication techniques, where cryptography-based authentication algorithms are commercial from 2G to 5G systems, blockchain-based authentication is a distributed framework-based group identity authentication approach, Physical-Layer Key-Based Authentication (PLKA) is achieved leveraging the reciprocity of wireless links between the transmitter and receiver, and Physical-Layer Authentication (PLA) is realized based on channel fingerprints and Radio Frequency (RF) fingerprints.
Techniques Descriptions
Cryptography-Based Authentication • 2G: Mainly applied in two systems: global systems for GSM and CDMA. The former generates signed response through key agreement protocols to realize access authentication, while the latter generates verification value through random number calculation. • 3G: Introducing the EAP-AKA protocol to realize the integrity protection of signaling, voice, and data. • 4G: Adopting the Evolved Packet System AKA (EPS-AKA) protocol, adopting signaling separation architectures of access and non-access layers, and resisting internal attacks through differentiated key distribution strategy, but lacking an effective key leakage response mechanism. • 5G: Realizing security upgrade based on EAP-AKA’ and 5G AKA protocols. Enhancements include introducing sequence numbers to defend against replay attacks and using asymmetric encryption to prevent international user identity information from being intercepted. •  Cons: Reliance on the limited computation power assumptions of attackers, difficulty in defense against replay attacks, and requirement for key generation, distribution and updates.
Blockchain-Based Authentication • distributed architecture-based group identity authentication method, with its characteristics primarily embodied in multi-tier frameworks. • During system operation, each node relies on consensus mechanisms and competitive computing to achieve continuous blockchain updates. When data from a single node is illegally tampered with, the system will automatically suspend the update permission of that node. Thanks to the fault-tolerant properties of distributed ledgers, communication between remaining nodes can still proceed normally. •  Constructing security barriers for user access through asymmetric encryption, while employing PKI to implement identity protections. •  Specifically, data layers ensure information immutability, network layers define terminal authentication mechanisms, consensus layers establish pre-agreed authentication protocols for all terminals, incentive layers motivate active participation in authentication tasks across all terminals, and contract layers contain intelligent identification algorithms and smart contracts for anomaly detection. •  Cons: High delay and low throughput.
PLKA • Be realized using physical-layer key, which is generated based on the inherent physical characteristics of wireless channels, including channel fading and noises. By measuring and analyzing these characteristics, keys closely related to specific links can be generated. • The reciprocity of channels ensures that the transmitter and receiver can generate consistent keys based on the same channel characteristics. • The physical characteristics are highly random and unpredictable, which endows the generated key with strong randomness and difficulty in copying, thus greatly improving the security of the key. • The physical characteristics of wireless channels change with time, environment, and other factors, which provides the possibility of frequently updating the key and further strengthens the security performance. •  The steps of its generation include channel detection, channel quantization, information negotiation, and privacy enhancement.It is highly sensitive to minor changes in input fingerprints. •  Cons: Difficult to implement in FDD systems and static environments.
PLA • Be realized based on physical-layer attributes, mainly including channel fingerprints and RF fingerprints. • Channel fingerprints are derived from the unique characteristics of signals affected by the propagation environment, such as channel state information (CSI), path loss, and delay spread. Channel fingerprints are unique in both time and space, and have short-term stability. • RF fingerprints originates from the manufacturing differences of equipment hardware, such as In-phased/Quadrature (I/Q) imbalance, carrier frequency offset (CFO), and phase noise. These differences lead to a unique “hardware fingerprint” embedded in the transmitted signal, which has hardware uniqueness and long-term stability. Even the devices of the same model have slight differences in their RF fingerprints, which are not easy to be covered up by software modification or environmental interference. •  Pros: Endogenous security. The endogenous security characteristics of channel fingerprints and RF fingerprints stem from the uniqueness and non-replicability of their physical properties, which are directly determined by the physical-layer characteristics of the communication system itself, without relying on external encryption algorithms or key management. • Pros: Lightweight. Firstly, it circumvents upper-layer signaling processes. Additionally, because access points have obtained CSI of legitimate users in channel estimation stages, the calculation overhead is further reduced. • Pros: High compatibility. Even if devices can not decode each other’s upper-layer signaling, as long as they share the same physical-layer communication standard, they can still decode the physical-layer bit stream and extract the fingerprints needed for PLA, thus realizing cross-device and cross-system authentication, which provides strong support for interoperability and compatibility between devices.
II-B GAI
II-B1 What is GAI?
GAI represents a category of AI technology aimed at synthesizing new data instances that adhere to a specified probability distribution. Unlike conventional DAI optimized for classification or regression tasks, GAI excels at learning potential structural patterns within training data, enabling autonomous generation of diverse content such as images, audio, and text. Its strength resides in modeling the distribution of high-dimensional data, facilitating the generation of realistic synthetic datasets, and augmenting sparse training samples. These capabilities make GAI particularly valuable in areas where data is scarce or labeling costs are too high, where DAI struggles.
II-B2 How to GAI?
Four typical models are presented as follows.
  • Generative Adversarial Network (GAN): GANs consist of a generator and a discriminator, participating in an adversarial two-player minimax game. The generator synthesizes data to simulate real samples, while the discriminator evaluates and distinguishes authenticity. Through iterative adversarial training, GANs excel at producing clear and high-fidelity outputs.
  • Variational Autoencoder (VAE): VAEs employ probabilistic encoder-decoder architectures to map data into a structured latent space and reconstruct it via probabilistic sampling. The loss function balances the reconstruction accuracy and KL divergence to align the latent distribution to a predefined Gaussian prior.
  • Generative Diffusion Model (GDM): GDMs learn to generate data by reversing a progressive noising process. By gradually denoising from random noise, they can model complex distributions with strong diversity and stability. Despite high computational complexity, they provide high-quality and consistent outputs.
  • Transformer: Transformers have become the cornerstone of Large Language Models (LLMs) through the unique self-attention mechanism and modular design. For example, ChatGPT combines supervised fine-tuning with reinforcement learning from human feedback to optimize conversational alignment, while DeepSeek incorporates a Mixture of Experts (MoE) architecture and introduces instruction tuning steps during training to improve efficiency and performance. These advancements enable LLMs to handle nuanced tasks, from creative writing to task-oriented dialogue.
TABLE II: The outline of the issues faced in the field of PLA, the limitations of Discriminative Artificial Intelligence (DAI) technologies in tackling these difficulties, and the promising capabilities of Generative AI (GAI) to boost performance within these contexts, where insufficient fingerprint data occur due to the unpredictable interference of environmental complexity and the inherent physical constraints of hardware measurement systems; environment noises and inferences appear due to path loss variations caused by obstacle obstruction, and rapid signal intensity fluctuations resulting from multipath effect; fingerprint data is perturbed by adversarial attacks and backdoor attacks; and complex tasks include multi-user authentication tasks, identifying new users joining the system, and realizing multiuser authentication and attack detection.
Issues Limitations of DAI Potential of GAI
Insufficient fingerprint data • It heavily relies on a large number of labeled fingerprint samples to define classification boundaries, and depends on upper-layer authentication protocols or manual operations to assign identity labels of fingerprint samples. • This strict supervised learning mode encounters constraints in annotation costs. When fingerprint samples are scarce, it is prone to overfitting or inadequate feature representation. • The potential value of unlabeled fingerprint samples remains underutilized, resulting in a waste of data resources. • It learns distributions of fingerprints rather than directly constructing mapping relationships between labels and fingerprints. • It not only significantly expands the size of training fingerprint sets but also allows for the controlled generation of fingerprint variants simulating different conditions, thereby enhancing the PLA model’s adaptability to data scarcity. • It takes the fingerprint data distribution itself as the learning object, rather than relying on identity labels, thus fully utilizing the potential value of unlabeled data [11].
Environment noises and inferences • It tends to treat noise as a static interference factor independent of fingerprint features, which restricts the model to constructing classification boundaries only under preset SNR conditions. • It prioritizes classification accuracy as its core optimization objective. The learning process often focuses on capturing the most easily distinguishable superficial features in noisy environments, but struggles to deeply analyze the complex nonlinear interactions between noise and fingerprints. • It constructs a joint probability space encompassing both fingerprints and noises. Thus, it explicitly expresses interaction through a mathematical structure, revealing the deep-seated patterns of noise’s intrinsic influence on fingerprint features [12]. • By incorporating SNR as a conditional input, It not only breaks the dependence of generated fingerprint samples on specific SNR conditions but also fundamentally improves the generalization when faced with varying SNRs [1].
Perturbations in fingerprint data • It is often limited to optimizing the classification boundaries of known clean fingerprint data, lacking proactive strategies to defend against adversarial attacks. • It typically relies on static rules or preset thresholds for authentication decisions, a rigid decision-making mechanism that is difficult to cope with evolving interference patterns. • It is highly sensitive to minor changes in input fingerprints. • GAN possesses an inherent defensive advantage due to its adversarial training nature [13]. • It cleverly formulates PLA as probability inference problems by learning the underlying probability distribution of fingerprints. • When input signals are disturbed, it can reconstruct the features of the original signal based on the learned fingerprint distribution, thereby eliminating the impact of perturbation components.
Complex tasks • When faced with highly similar or significantly overlapping fingerprint features among multiple users, the classification boundaries are forced to be compressed into narrow regions, leading to ambiguity in decision-making. • It requires re-adjusting classification boundaries to adapt to the updated user categories. • It typically requires separately constructing multiuser authentication models and attack detection models, and achieving collaboration between them through series or parallel connections. • It builds an exclusive probability distribution model for the fingerprint features of each user individually, rather than forcing fingerprint features of all users into the same space. • It only needs to train the generator branch for updated users without modifying existing user modules or the global model structure. This “plug-and-play” feature greatly reduces the maintenance complexity of the PLA system. • It integrates multiuser authentication and attack detection into a single-stage task through joint probability modeling [5].
II-C What PLA Issues Can Benefit from GAI?
As depicted in Table II, the limitations of DAI and GAI’s potential in PLA are presented as follows.
  • Insufficient Fingerprint Data: Given the unpredictable interference caused by environmental complexity and the inherent physical constraints of hardware measurement systems, it is often difficult to collect sufficient high-precision fingerprint samples in actual communication scenarios [3]. Traditional DAI methods have limitations in relying on a large number of fingerprint samples, relying on upper-layer authentication protocols or manual labeling of fingerprints, and low utilization of unlabeled fingerprint data. In contrast, GAI has potential in augmenting fingerprint data, learning fingerprint data distributions through latent variables, learning unlabeled fingerprint data, and continuously optimizing its model parameters and structure.
  • Environment Noises and Inferences: In wireless environments, the signal-to-noise ratio (SNR) often fluctuates due to various factors, including changes in path loss caused by mobile device displacement, shadow fading induced by obstacle obstruction, and rapid signal intensity fluctuations resulting from multipath effects [1]. Traditional DAI methods have limitations in training PLA models for specific SNR conditions, capturing the coupling relationship between fingerprints and noises or inferences, and requiring prior knowledge of noises. In contrast, GAI has potential in modeling the joint distribution of fingerprints and noises, synthesizing fingerprints under different SNRs through conditional generation, and denoising and reconstructing fingerprints.
  • Perturbations in Fingerprint Data: Faced with emerging intelligent threats such as adversarial attacks [13] and backdoor attacks [2], traditional DAI methods have limitations, lacking defense against well-designed adversarial perturbations, decision making based on fixed rules, and vulnerability and sensitivity to small disturbances. In contrast, GAI has the potential to defend against adversarial attacks through adversarial training, eliminate interference through fingerprint reconstruction, and make flexible inferences, judgments, and decisions based on learned fingerprint probability distributions.
  • Complex Tasks: For complex tasks such as multiuser authentication and identifying updated user categories, traditional DAI methods have limitations in fuzzy classification boundaries caused by overlapping fingerprint feature spaces of users and require retraining the PLA model to adapt to the updated user categories. In contrast, GAI has the potential to generate independent fingerprint distributions for each user to avoid feature overlap and expand to updated users without retraining the PLA model.
Figure 1: Illustration of GAI-enhanced methods for PLA, where GAI models include Generative Adversarial Network (GAN), Variational Autoencoder (VAE), Generative Diffusion Model (GDM), and Transformer; and PLA issues are classified according to fingerprint collection, model training, and performance optimization phases. The fingerprint collection stage provides fingerprint dataset for model training stage, and the performance of the trained PLA model is further enhanced in the performance optimization stage.
TABLE III: The use of GAI in PLA.
Phases Issues GAI Models
GAN VAE GDM Transformer
Fingerprint Collection Fingerprint Estimation Determining mapping functions of fingerprints through generators [10] Approximating estimation through variational inference process Estimating fingerprints through posterior sampling -
Fingerprint Augmentation Conditioning GANs for time-varying channel environments [6] Decoders for augmenting fingerprints Generating fingerprint samples during reverse processes [12] -
Fingerprint Filling - - - Learning dependencies between fingerprints by attention mechanisms [11]
Fingerprint Transmission - Encoders and decoders for joint source-channel coding Channel adaptive de-noising transmission -
Model Training Attack Detection Modifying the discriminator’s functionality to output category distributions Latent Space for unsupervised classification [3] Comparing generation probabilities under different identity hypotheses -
Multiuser Authentication Formulating as an OSR problem [5] - - -
Efficient Training and Deployment Generating diverse fingerprints to meet the requirements of complex environments Supporting probabilistic modeling and feature decoupling Formulating as training a conditional denoiser [4] Supporting complex logical reasoning and knowledge transfer
Performance Optimization Flexibility for Eliminated Legitimate Users Gradient-based optimization Gradient-based optimization Gradient-based optimization, data sharding, and adding learnable layers Gradient-based optimization, knowledge distillation, and adding learnable layers
Defense Against Intelligent Attacks Adversarial training [13] Identifying abnormal samples through reconstruction errors Eliminating fingerprint triggers through denoising process -
Adaptation for Different Scenarios - - - Parameter-efficient fine-tuning techniques [9]
III GAI-enhanced Methods for PLA
As illustrated in Figure 1, the implementation of GAI-based PLA necessitates a self-evolving system comprising three stages: fingerprint collection, model training, and performance optimization.
  • During the fingerprint collection phase, GAI empowers fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission to improve accuracy, enrich diversity, ensure completeness, and enhance transmission efficiency, respectively, thereby providing reliable fingerprint samples for training PLA models.
  • In the model training phase, GAI-driven training strategies are tailored to specific task requirements, such as unsupervised attack detection, multi-user authentication, and efficient training and deployment, to develop PLA models.
  • During the performance optimization phase, the trained PLA model undergoes further enhancement through GAI to achieve advanced capabilities, including improving flexibility for eliminated legitimate users, enhancing robust defense against intelligent attacks, and boosting adaptability across diverse scenarios.
Table III further compares how representative GAI models address PLA issues from each of these phases, with detailed descriptions provided below.
III-A Fingerprint Collection
III-A1 Fingerprint Estimation
Channel estimation is a critical technique for acquiring channel fingerprints. DAI-based estimation methods require long pilot sequences to achieve satisfactory performance and struggle in high path loss scenarios. In contrast, GAI can generate synthetic training fingerprints with distributions similar to real training samples, thereby providing effective support for fingerprint estimation models. For instance, the variational inference process in VAEs approximates the maximum likelihood process of fingerprint estimation. Similarly, GDMs can learn score functions of posterior fingerprint distributions and employ posterior sampling during reverse processes to recover fingerprint samples. Additionally, Zhang et al. [10] applied the discriminator in GANs to learn and extract intrinsic features of time-varying channels in high-speed railway systems, while the generator determines implicit mapping functions of training fingerprint data to enhance estimation performance.
III-A2 Fingerprint Augmentation
GAI can augment fingerprint datasets, thereby enhancing the robustness of PLA models. For instance, Jiang et al. [6] employed conditional GANs to synthesize fingerprint samples for Low Earth Orbit (LEO) satellites. This conditional approach enables the synthesis of more realistic fingerprints, as environmental factors are explicitly incorporated into the generation process. Similarly, Chi et al. [12] employed GDMs to jointly leverage temporal, spectral, and complex-valued features of fingerprint waveforms, thus enhancing the generation performance of fingerprints.
III-A3 Fingerprint Filling
In real-world, fingerprint data loss may occur due to physical obstructions, device malfunctions, and environmental disturbances. Such missing fingerprint data directly disrupts the spatiotemporal continuity of fingerprint sequences, necessitating fingerprint filling methods to restore complete fingerprint sequences. DAI-based methods establish data correlations only between adjacent nodes through local neighborhood modeling, making them ill-suited to adapt to dynamic changes in node positions or obstruction patterns. In contrast, GAI models, such as Transformers, leverage self-attention mechanisms to automatically learn dependency strengths between different fingerprint data points, dynamically capturing long-range spatiotemporal correlations, thereby achieving more robust missing fingerprint reconstruction performance. For example, Zhao et al. [11] introduced a Bidirectional Encoder Representations from Transformers (BERT)-based framework for robust fingerprint filling.
III-A4 Fingerprint Transmission
Compared with centralized PLA, collaborative PLA utilizing multi-party cooperation exhibits superior authentication performance. As the number of dimensions of fingerprint samples increases, a key challenge is how to efficiently transmit the fingerprints estimated by collaborators to Bob. Semantic communication uses joint source-channel coding to understand the semantic content of information rather than just focusing on accurate symbol transmission, thereby reducing bandwidth requirements and potentially improving the efficiency and reliability of fingerprint transmission. GAI can help further enhance the performance of fingerprint transmission based on semantic communication, such as the construction of fingerprint knowledge bases and the management of fingerprint transmission models.
III-B Model Training
III-B1 Attack Detection
GAI, leveraging self-supervised learning mechanisms, models the probability distribution of fingerprints without relying on manually annotated “fingerprint-identity” mapping relationships. Although it was not originally designed for classification tasks, through architectural adjustments, it remains capable of performing unsupervised attack detection. For example, Meng et al. [3] combined the advantages of AEs and VAEs to obtain superior attack detection performance when processing fine-grained, high-dimensional fingerprint samples. Additionally, by modifying the discriminator’s functionality to output category probability distributions while distinguishing data authenticity, unsupervised classification can be achieved using GANs. Moreover, by comparing generation probabilities under different identity category hypotheses, GDM can implicitly construct detection boundaries without explicit identity labels.
III-B2 Multiuser Authentication
In communication systems, PLA models need to authenticate multiple legitimate users simultaneously and detect unknown attackers. Open-Set Recognition (OSR) provides an effective solution to this challenge. OSR requires the model to not only correctly classify known classes seen during training, but also effectively recognize and reject samples of unknown classes during the testing phase. Compared with traditional OSR methods, such as classifier-based adjustment approaches and distance metric methods, GAI has the following advantages: excelling in detecting high-dimensional unknown fingerprint samples, offering more direct detection evidence through reconstruction errors, and demonstrating stronger adaptability to unknown fingerprint categories. For instance, Guo et al. [5] introduced GANs to distinguish between legitimate and anomalous devices.
III-B3 Efficient Training and Deployment
Channel Knowledge Map (CKM) helps communication systems predict CSI and reduce the overhead of real-time fingerprint estimation by constructing a location-specific CSI database. By building CKM, CSI-based downstream tasks such as PLA can be effectively empowered. Compared with traditional CKM construction methods, GAI can reduce dependence on measured CSI data, generate CSI fingerprints for specific locations, and achieve efficient transfer of different channel models. For instance, Fu et al. [4] treated incomplete CKM images as conditions for training a conditional denoiser within the decoupled GDM, thereby addressing the pivotal challenge of reconstructing a comprehensive map for all relevant locations using only partially observed channel knowledge data.
III-C Performance Optimization
III-C1 Flexibility for Eliminated Legitimate Users
As the departure of old users, GAI-based PLA models need to forget the fingerprint characteristics of old users or incorporate the relevant distributions of new users to optimize and update authentication parameters. Four machine unlearning strategies can be applied to forget the fingerprint characteristics of old users, including gradient-based optimization, knowledge distillation, data sharding, and adding learnable layers strategies.
III-C2 Defense Against Intelligent Attacks
Attackers can generate adversarial samples to deviate the fingerprint features extracted at the receiver from the true distribution, thereby deceiving PLA models [13]. Additionally, attackers can embed fingerprint data with specific trigger patterns before the training phase, and activate this backdoor during the authentication phase [2]. Papangelo et al. [13] verified that adversarial training can defend against adversarial attacks for PLA models. Additionally, adversarial training can also defend against backdoor attacks using the discriminator. Moreover, VAEs can identify abnormal samples through reconstruction errors. Besides, the denoising process of GDM models can be used to eliminate fingerprint triggers, disrupting the structural consistency of the triggers.
III-C3 Adaptation for Different Scenarios
GAI, especially Transformer-based LLMs, with its powerful generalization ability and multi-scenario knowledge transfer, can flexibly adapt to the requirements of PLA in different application scenarios. In the pre-training phase, LLM learns fingerprint feature distributions from various scenarios. When encountering new scenarios, LLM adopts parameter-efficient fine-tuning techniques to adapt to local fingerprint features by updating some parameters, thereby avoiding the computational cost of retraining the entire model. For instance, Liu et al. [9] combined LLMs and proposed a BERT-based lightweight PLA framework to enhance the zero-trust edge IoT security.
IV Case Study: GAI-Based PLA for Dynamic Environments
Figure 2: Illustration of the proposed GDM-based PLA scheme for dynamic environments. During offline training, Alice’s fingerprints $X_{A}$ and Jack’s fingerprints $X_{J}$ are jointly pre-processed as one image $X$, and GDM is utilized to learn the joint fingerprint distributions. During online authentication, the trained GDM is leveraged to output predicted Alice’s fingerprints $\hat{X}_{A}$ according to Jack’s fingerprints, and the identities of estimated unknown fingerprints $\tilde{X}$ are obtained by comparing them with $\hat{X}_{A}$.
IV-A Proposed GDM-based PLA Scheme
IV-A1 Challenges
Existing PLA schemes based on CSI fingerprints face challenges in dynamic wireless environments, where sudden environmental fluctuations cause drastic changes in the spatial distribution and temporal dependence of CSI fingerprints. These rapid changes make the previously learned decision boundaries no longer applicable to the current channel environment, thereby constraining authentication performance.
IV-A2 Motivations
Based on channel extrapolation principles [14], we hypothesize correlations between the CSI fingerprints of the legitimate transmitter (Alice) and collaborator (Jack). We further employ GDM to learn the association between their CSI fingerprints. During authentication, by leveraging Jack’s fingerprints as the condition, the trained GDM can improve the prediction accuracy of Alice’s fingerprints, thus enhancing the authentication performance.
IV-A3 System Model
As shown in Figure 2, Alice is a legitimate device that requires identity authentication performed at Bob, Jack is a collaborative device, and Eve is a spoofing attacker impersonating Alice. We assume that Alice and Jack are both stationary, while the surrounding environment changes rapidly. The fixed and close distance between Alice and Jack allows for the establishment of a stable relationship between their CSI fingerprints through channel extrapolation in the user domain [14].
IV-A4 Steps of the Proposed Scheme
As illustrated in Figure 2, the proposed scheme includes the following steps:
  • Offline Training: Bob obtains Alice’s fingerprints $X_{A}$ and Jack’s fingerprints $X_{J}$ through channel estimation. Their distributions are further learned through GDM. The problem is formulated as fitting a function $X_{A}=f(X_{J})$ to predict $X_{A}$ using $X_{J}$. Here, we train GDM to represent $f$. Specifically, it gradually adds noise to disrupt fingerprint data, and then learns to recover the original data from the noise through a reverse denoising process. The loss function of the denoising process is $\mathcal{L}=\mathbb{E}_{X_{A},X_{J},\epsilon_{t},t}\left\|\epsilon_{t}\;-\;\epsilon_{\Theta}\bigl{(}X_{A,t},\,t,\,X_{J}\bigr{)}\right\|_{2}^{2}$, where $\epsilon_{t}$ is the Gaussian noise added to $X_{A}$ at diffusion step $t$, $X_{A,t}$ is the noisy fingerprint after $t$ diffusion steps, and $\epsilon_{\Theta}(X_{A,t},t,X_{J})$ is the predicted noise by model parameters $\Theta$.
  • Online Authentication: Bob obtains Jacks’ fingerprints $X_{J}$, and leverages the trained GDM to predict Alice’s fingerprint $\hat{X}_{A}$. Then, Bob estimates fingerprints $\tilde{X}$ from unknown transmitters, and predict the identities by comparing the differences between $\hat{X}_{A}$ and $\tilde{X}$.
Figure 3: Simulation parameters and simulation results. During training, we have 10, 860 fingerprint samples for Alice and Jack. During authentication, we run 50 times to obtain the average result, and we have 50 fingerprints for Alice, Eve, and Jack in each time. We use $F_{1}=(2P_{tl}P_{ta})/(P_{tl}+P_{ta})$ and authentication error rate $R_{e}$ to evaluate the authentication performance [3]. $F_{1}\in[0,1]$, and larger $F_{1}$ means better authentication performance. $P_{tl}$ and $P_{ta}$ respectively represent correct authentication rate and false alarm rate. $R_{e}\in[0,1]$, and larger $R_{e}$ means worse authentication performance. The proposed scheme utilizes GDMs to generate Alice’s fingerprint at time $T$ based on Jack’s fingerprint at time $T$, and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 1: The fingerprint to be authenticated at time $T$ is compared with Alice’s fingerprint at time T1T-1. Baseline Scheme 2: The typical time series model, Long Short-Term Memory (LSTM) model, is used to predict Alice’s fingerprint at time $T$ and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 3: The fingerprint to be authenticated at time $T$ is compared with Jack’s fingerprint at time $T$.
IV-A5 Complexity Analysis
The computational complexity is denoted as $\mathcal{O}\left(TBLN\left(2C^{2}{S}+4{S}^{2}C\right)\right)$, where $T$ is the total time step, $B$ is the batch size, $L$ is the number of layers, $N$ is the number of ResNetBlocks in each layer, and $C$ and S{S} denote the maximum channel number and spatial resolution, respectively. $2$ reflects the dual-branch design, while $4$ in the attention term arises from the addition of cross-attention alongside self-attention in every block. According to the simulation results on the computer configured as an NVIDIA RTX A6000 GPU paired with an Intel Xeon w7‑2495X CPU, each time step requires 18.05 ms.
IV-B Simulation Results
The DeepMIMO datasets are employed to generate CSI fingerprint samples for simulation experiments. To preserve the phase integrity of complex fingerprints, we compute the dynamic ranges of its real and imaginary components as the differences between their maxima and minima, and then use the larger of these two ranges as a common normalization factor to linearly map both the real and imaginary values into the interval [1,1][-1,1]. Figure 3 illustrates the DeepMIMO parameters, GDM parameters, the training process, and the simulation results. The proposed scheme can accurately recognize the identity of each fingerprint during authentication, but the three baseline schemes can not. Due to the time-varying nature of wireless channels and the continuous change in Alice’s position, there are deviations in the distribution of Alice’s channel fingerprints between the training and authentication phases. The proposed scheme can learn the relative features between Alice’s and Jack’s fingerprints through GAI and predict Alice’s fingerprints based on Jack’s fingerprints during authentication, thereby improving authentication performance in dynamic environments. In contrast, baseline scheme 1 only utilizes Alice’s fingerprints from the previous moment, which differs from the current distribution of Alice’s fingerprints. Although baseline scheme 2 employs LSTM to learn the distribution changes of fingerprints, its ability to learn fingerprint distributions is weaker than that of GAI. Because there are correlations between Alice’s and Jack’s fingerprint distributions, baseline scheme 3 can obtain better authentication performance in low SNR environments than baseline schemes 1 and 2.
IV-C Lessons Learned
The case study leverages channel extrapolation principles and GDMs to model and learn the joint distributions of Alice’s and collaborator’s channel fingerprints. Due to the excellent generation ability and noise robustness of GDMs, the authentication performance of this scheme in dynamic environment is significantly better than that of the three baseline schemes.
V Challenges and Future Directions
V-A Challenges
V-A1 Computational Demands and Latency Implications
Although GAI can enhance the performance of PLA, its high computing power requirements also bring some challenges to its deployment. Currently, the growth of GAI model sizes has outpaced advancements in general-purpose compute hardware. Both parameter counts and per-request floating-point operations (FLOPs) are scaling more rapidly than accelerator performance improvements, creating escalating compute footprints for both training and inference. Additionally, the inference time of GAI-based PLA models will have a significant impact on the communication latency. Especially in the dynamic environment, with the continuous change of fingerprint identity characteristics, high-frequency identity authentication is required to ensure the legitimacy of identity.
V-A2 Security of GAI-based PLA Models
During the training process of GAI model, it directly contacts and remembers the features of fingerprint data. Intelligent attackers may launch member inference attacks to determine whether a specific fingerprint data exists in the training dataset. They could further reverse engineer privacy sensitive fingerprint data such as the location of legitimate users. Additionally, attackers may tamper with the training fingerprint data of GAI models through adding adversarial data, causing the GAI model to learn incorrect fingerprint distributions. Attackers may also implant backdoors during training, which hide a mapping relationship between triggers and malicious outputs. When encountering specific triggering patterns, these backdoors may trigger pre-defined error outputs.
V-A3 Trade-off Issues
The trade-off between security performance and complexity in GAI-based PLA is essentially about improving security through GAI technology while controlling the computational, data, and hardware overhead introduced as a result. As mentioned above, GAI can enhance the fingerprint collection, model training, and performance optimization stages of PLA. Although GAI technology improves access security, it brings challenges such as computational complexity and hardware requirements, which increase deployment costs. In addition, GAI may introduce new security threats, such as adversarial attacks, which also require sacrificing complexity to further enhance the reliability of GAI models.
V-A4 GAI-based PLA for Emerging Scenarios
As a typical scenario for 6G, ISAC achieves the goal of spectrum conservation and mutual enhancement between communication and sensing. However, there are inconsistencies in the spatiotemporal dimensions of channel fingerprints, RF sensor fingerprints, and non-RF sensor data. Additionally, to support secure access for massive communication in 6G, PLA models need to pre-register and learn the fingerprint features of a massive number of devices. Besides, 6G will also support Hyper Reliable and Low-Latency Communication (HRLLC). However, GAI models typically require high computational resources. Thus, the inference time may exceed the latency budget, especially in resource constrained IoT devices, which cannot meet the requirements of real-time PLA.
V-B Future Directions
V-B1 Model Compression for GAI-based PLA
Model compression can not only reduce the complexity of GAI models, but also potentially defend against new security threats such as backdoor attacks. Model compression techniques mainly include model pruning, parameter quantization, low-rank decomposition, and knowledge distillation. For example, knowledge distillation transfers the knowledge of the teacher model to the student model while preserving core features, and model pruning can reduce erroneous outputs in specific input fingerprint patterns by removing redundant neurons or connections, thus removing implanted backdoor triggering features.
V-B2 Federated Learning for GAI-based PLA
Federated learning [15] provides a foundational framework for trusted training of GAI-based PLA models through a distributed training architecture. It allows legitimate users to collaborate in training global GAI models without sharing locally collected fingerprint datasets. It can be combined with differential privacy to further enhance the privacy protection capability of GAI-based PLA models.
V-B3 Distributed Trust for GAI-based PLA
Distributed trust builds a decentralized trust mechanism to ensure the credibility of the behavior of each distributed node and the overall security of the system. By using distributed ledger technology to record the source of training fingerprint data for GAI, it ensures that the fingerprint data is traceable and tamper proof. Additionally, multi-party secure computing (MPC) or trusted execution environment (TEE) can be combined to process fingerprint samples in a distributed environment, protecting user privacy.
V-B4 Emerging GAI Models for PLA
Some emerging GAI models, such as Large Language Models (LLMs), can be introduced to further enhance the performance of PLA. By utilizing the powerful representation and generation potential of LLMs, an efficient and dynamically evolving fingerprint database can be constructed. LLMs can generate fingerprints in different physical environments, covering extreme scenarios and solving the problem of scarce fingerprint data. In addition, by inputting environmental parameters, LLM can generate fingerprints that are suitable for the current environment. Moreover, LLMs can be used to generate semantic labels for fingerprints to support fast retrieval and matching of fingerprints.
V-B5 GAI-based PLA for 6G
GAI plays a critical role in learning multi-modal fingerprints for ISAC. It constructs a unified cross-modal representation space through latent space mapping, effectively resolving the alignment challenges posed by heterogeneous data. Additionally, in scenarios where partial sensors fail, the cross-modal reasoning capabilities of GAI models enable the reconstruction of consistent representations for missing modalities, ensuring the completeness of the authentication system’s input. Additionally, LLMs are promising to support secure access in massive communication and ubiquitous connectivity. By fine-tuning the pre-trained LLMs, the customized authentication needs of specific scenarios can be met.
VI Conclusions
In this paper, we have first revisited classic identity authentication techniques, including cryptography-based authentication, blockchain-based authentication, PLKA, and PLA. Then, we have reviewed representative GAI models, including GANs, VAEs, GDMs, and Transformers, and we have analyzed the limitations of applying DAI in PLA and have pointed out the potential of GAI-based PLA. Furthermore, we have provided GAI-enhance methods for PLA in detail, including fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission in fingerprint collection phase, attack detection, multiuser authentication, and efficient training and deployment in model training phase, and flexibility for eliminated legitimate users, defense against intelligent attacks, and adaptation to different scenarios in performance optimization phase. Additionally, we have presented a case study to show the superiority of GAI over DAI in authentication performance under dynamic environments. Lastly, we have provided challenges and future directions.